The standard deviation of a binomial distribution for which n = 100 and p = 0.40 is 24.0.
True or False?

Answer by Edwin McCravy(20081)

The formula for the standard deviation of a binomial distribution is

s = √npq

where q = 1-p = 1-0.40 = 0.60

or you may have learned the formula without q, but with 1-p like this:

s = √np(1-p)

s = √100(0.40)(0.60)

s = √24 = 4.8990

So it's the square root of 24, (not 24 without a square root over it.)

Answer: false

Note: the formula for the mean is m = np without a square root.
But the formula for the standard deviation does have a square root.
If you had been asked for the variance s² = npq, not the standard 
deviation, then 24 would have been the answer, but the standard 
deviation is the square root of the variance.

Edwin